Circular No. 5615

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA IN NGC 1313? S. D. Ryder and M. A. Dopita, Mount Stromlo and Siding Spring

Observatories (MSSSO); L. Staveley-Smith, Australia Telescope National Facility, CSIRO; D. Malin, Anglo-Australian Observatory; E. Colbert, University of Maryland; R. Petre and E. Schlegel, Goddard Space Flight Center; and D. Campbell-Wilson, University of Sydney, report: “The suspected nova in NGC 1313 (IAUC 4950) is possibly a supernova with characteristics similar to SN 1961V and SN 1986J (type IIpec or V). Archival plates reveal that an outburst occurred in mid-1978, the brightest known image being that in Fig. 3 of Marcelin and Athanassoula (1982, A.Ap. 105, 76); the nova/supernova candidate is 15 mm to the left and 8 mm up from the lower right corner of this reproduced plate, or ∼ 218′′ west and 193′′ south of the optical center of the galaxy quoted in the ESO/Uppsala Atlas. Approximate magnitudes, from U.K. Schmidt Telescope plates unless noted otherwise (SB = Sandage and Bedke 1988, Atlas of Galaxies, NASA SP-495; ES = ESO Schmidt plate; S = CCD photometry by Ryder using the MSSSO 1.0-m and 2.3-m telescopes): 1973 Oct. 1 and Nov. 24, B > 22 (ES); 1974 Sept. 16, 18, Oct. 12, 1975 Jan. 11, Oct. 3, 28, and Nov. 29, B_{J} = 22.1; 1976 Dec. 14, R = 20.0; 1977 Oct. 12, B_{J} > 22.1; 1978 July 31, B = 16.0 (Marcelin and Athanassoula; see above); Sept. 26, B = 17.9 (SB; panel 51); Oct. 4, V = 17.7 (SB; panel 7); Nov. 24, B_{J} = 19.0; 1981 Oct. 7, B_{J} = 22.1; 1982 Nov. 14, B_{J} > 22.1; 1984 Sept. 27, R = 20.0 (ES); 1986 Oct. 23, R = 18.4 (ES); Nov. 22, B_{J} = 21.7; 1987 Oct. 26, R = 18.4; Oct. 27, B_{J} = 21.6; Nov. 25 and Dec. 24, I > 20.0; 1988 Jan. 11, R = 18.4; Dec. 14, U > 21.0, B > 20.7; 1989 Nov. 22, R = 18.1; Dec. 29, V = 20.0 (S); 1990 Oct. 24, R = 18.0; 1992 Jan. 5, V = 20.30, R = 18.70, B = 20.90 (S). Observations in 1992 July with the Australia Telescope Compact Array show a strong non-thermal radio source (147 mJy at 1.4 GHz) located at α = 3^{h}17^{m}02^{s}.4, δ = −66^{o}43′57′′ (equinox 1950.0), within 1′′ of the optical position. Observations of NGC 1313 at 843 MHz with the Molonglo Observatory Synthesis Telescope in 1982 January (Harnett 1987, MNRAS, 227, 887), 1986 January, and 1992 July give respective flux densities of 90, 224, and 157 mJy, consistent with a powerful radio supernova in a dense circumstellar environment. ROSAT observations in 1991 April and May reveal an extremely powerful x-ray source with an unabsorbed flux of 3.4 × 10^{−12} erg s^{−1} cm^{−2} at the same position. This source was not present in an Einstein image taken in 1980 January, again consistent with a supernova outburst in 1978.”

1992 September 22 Daniel W. E. Green